# Packing Input Frame Context in Next-Frame Prediction Models for Video Generation

### v2

@@ 4.5 Comparison to Alternative Architectures
Noisy history: Delay the denoising timestep on history latents so that history latents are noisy (but less noisy than the current generating latents). Reducing the reliance on the history is beneficial for interrupting error accumulation, thus mitigates drifting, but at the cost of aggravating forgetting. We implement a structure that resembles DiffusionForcing [4].
History guidance: Delay the denoising timestep on history latents but also put the completely noised history on the unconditional side of CFG guidance. This will speed up error accumulation thus aggravating drifting, but also enhance memory to mitigate forgetting. We implement a structure that resembles HistoryGuidance [34].
Table 2: Comparison with relevant architectures. We compare with other relevant methods across the same global metrics, drifting metrics, and human assessments used in the ablation study. The tests are conducted with HunyuanVideo as base. Bests in bold. ELO differences within ±$\pm 16$ are considered ties. Method Global Metrics $\uparrow$ Drifting Metrics $\downarrow$ Human Clarity Aesthetic Motion Dynamic Semantic Anatomy Identity $\Delta_{\text{drift}}^{\text{Clarity}}$ $\Delta_{\text{drift}}^{\text{Motion}}$ $\Delta_{\text{drift}}^{\text{Semantic}}$ $\Delta_{\text{drift}}^{\text{Anatomy}}$ ELO $\uparrow$ Rank $\downarrow$ Repeating image-to-video 56.73% 56.15% 94.34% 65.56% 17.74% 69.41% 73.06% 9.51% 3.92% 9.95% 19.88% 1016 5 Anchor frames (resembling StreamingT2V [15]) 69.58% 67.35% 99.96% 74.97% 25.76% 85.01% 79.52% 2.85% 2.15% 3.45% 9.25% 1172 2 Causal attention (resembling CausVid [53]) 62.88% 59.41% 96.98% 67.56% 19.15% 72.74% 75.32% 7.45% 3.15% 6.75% 15.96% 1087 4 Noisy history (resembling DiffusionForcing [4]) 66.75% 63.78% 97.20% 91.45% 22.98% 77.60% 77.92% 4.35% 2.05% 3.15% 9.85% 1146 3 History guidance (resembling HistoryGuidance [34]) 68.05% 68.74% 97.01% 73.39% 24.88% 81.84% 83.42% 7.35% 2.21% 5.25% 12.78% 1151 3 Proposed (f1k1_x_g9_f1k1f2k2f16k4) 71.15% 68.71% 99.45% 89.29% 28.15% 86.53% 82.11% 2.25% 1.85% 2.68% 8.58% 1221 1
As shown in Table 2, we observe several findings: (1) The proposed method achieves the best results in 3 global metrics, while other methods excel in at most one or two metrics. (2) The proposed method achieves the best results across all drifting metrics. (3) This evaluation aligns with human perceptions as evidenced by the ELO score.
@@ 5 Conclusion
In this paper, we presented FramePack, a neural network structure that aims to address the forgetting-drifting dilemma in next-frame prediction models for video generation. FramePack applies progressive compression to input frames based on their importance, ensuring the total context length converges to a fixed upper bound regardless of video duration. This is achieved through manipulating transformer patchify kernel sizes for different compression rates. Combined with the anti-drifting sampling methods that incorporate bi-directional context through early-established endpoints or inverted temporal ordering, our approach enables longer video generation while maintaining unchanged computational bottlenecks. Experiments suggest that FramePack can improve model responsiveness and visual quality in inference, while allowing for higher batch sizes in training. The approach is compatible with existing video diffusion models and supports various compression variants that can be optimized for wider applications.

### v3

@@ 5.4 Comparison to Alternative Architectures
Table 2: Comparison with relevant methods. We compare across the same global metrics, drifting metrics, and human assessments. The tests are conducted with HunyuanVideo as base. Bests in bold. ELO differences within $\pm 16$ are considered ties. Method Global Metrics $\uparrow$ Drifting Metrics $\downarrow$ Human Clarity Aesthetic Motion Dynamic Semantic Anatomy Identity $\Delta_{\text{drift}}^{\text{Clarity}}$ $\Delta_{\text{drift}}^{\text{Motion}}$ $\Delta_{\text{drift}}^{\text{Semantic}}$ $\Delta_{\text{drift}}^{\text{Anatomy}}$ ELO $\uparrow$ Rank $\downarrow$ Repeating image-to-video 56.73% 56.15% 94.34% 91.21% 17.74% 69.41% 73.06% 9.51% 3.92% 9.95% 19.88% 1015 5 Anchor frames (resembling StreamingT2V [20]) 69.58% 67.35% 99.96% 74.97% 25.76% 85.01% 79.52% 2.85% 2.15% 3.45% 9.25% 1173 2 Causal attention (resembling CausVid [65]) 62.88% 59.41% 96.98% 88.27% 19.15% 72.74% 75.32% 7.45% 3.15% 6.75% 15.96% 1087 4 DiffusionForcing [6] ($\sigma_{\text{train}}$ random, $\sigma_{\text{test}}=0.1$) 66.08% 65.76% 96.32% 91.59% 23.14% 75.93% 74.47% 4.84% 2.54% 3.33% 10.99% 1170 2 DiffusionForcing [6] ($\sigma_{\text{train}}$ random, $\sigma_{\text{test}}=0.5$) 67.41% 66.66% 92.93% 91.08% 24.03% 77.83% 76.42% 3.55% 2.39% 3.48% 9.40% 1174 2 DiffusionForcing [6] ($\sigma_{\text{train}}$ random, $\sigma_{\text{test}}=0$) 66.99% 64.73% 96.47% 90.45% 21.09% 80.62% 78.77% 8.41% 3.80% 8.47% 17.44% 1095 4 DiffusionForcing [6] ($\sigma_{\text{train}}=0.1$, $\sigma_{\text{test}}=0.1$) 66.19% 68.60% 94.89% 91.89% 22.49% 76.12% 78.27% 6.82% 3.79% 5.08% 10.78% 1149 3 History guidance (resembling HistoryGuidance [41]) 68.05% 68.74% 97.01% 73.39% 24.88% 81.84% 83.42% 7.35% 2.21% 5.25% 12.78% 1152 3 Inverted anti-drifting (f1k1_x_g9_f1k1f2k2f16k4) 71.15% 68.71% 99.45% 89.29% 28.15% 86.53% 82.11% 2.25% 1.85% 2.68% 8.58% 1220 1 Vanilla + discrete history (f16k4f2k2f1k1_g9+D, $K=256$) 70.01% 68.76% 95.65% 91.74% 28.37% 86.41% 82.22% 3.13% 2.05% 2.89% 8.74% 1224 1
@@
DiffusionForcing: We conduct a detailed ablation study with DiffusionForcing [6]. We focus on the history noise scheduling using the same scheduling as SkyreelV2 [7] that multiplies the diffusion timestep on history frames with $\sigma_{\text{train}}$ during training and $\sigma_{\text{test}}$ in inference to delay the denoising on history latents. Intuitively, $\sigma_{\text{test}}=0$ is equivalent to clean latents for history (no noise added to the history). We consider these ablations: (1) $\sigma_{\text{train}}$ being random with $\sigma_{\text{test}}=0.1$; (2) $\sigma_{\text{train}}$ being random with $\sigma_{\text{test}}=0.5$; (3) $\sigma_{\text{train}}$ being random with $\sigma_{\text{test}}=0$; (4) $\sigma_{\text{train}}=0.1$ and $\sigma_{\text{test}}=0.1$. Usually higher $\sigma_{\text{test}}$ reduces the reliance on the history, which is beneficial for interrupting error accumulation, thus mitigates drifting, but at the cost of aggravating forgetting.
History guidance: Delay the denoising timestep on history latents but also put the completely noised history on the unconditional side of CFG guidance. This will speed up error accumulation, thus aggravating drifting, but also enhance memory to mitigate forgetting. We implement a structure that resembles HistoryGuidance [41].
As shown in Table 2, we observe several findings. (1) The inverted anti-drifting sampling achieves the best results across all drifting metrics, while having a relatively small dynamic range. (2) The vanilla sampling with discrete history is very competitive in drifting measurements, while having a relatively larger dynamic range. (3) Human perception prefers the two proposed candidates as evidenced by the ELO score. (4) See also the supplementary material for more detailed explanations, analysis, and comparisons with DiffusionForcing candidates.
@@ 6 Conclusion
In this paper, we presented FramePack, a neural network structure that aims to address the forgetting-drifting dilemma in next-frame prediction models for video generation. FramePack applies progressive compression to input frames based on their importance, ensuring the context length converges to a fixed upper bound. We discussed both time-proximity-based packing and feature-similarity-based packing. We also discussed the anti-drifting training methods with history discretization, and the anti-drifting sampling methods using bi-directional context planning and scheduling. Experiments suggest that FramePack can process a large number of frames, improve model responsiveness, and allow for higher batch sizes in training. The approach is compatible with existing video diffusion models and supports various compression variants that can be optimized for wider applications.
Acknowledgments and Disclosure of Funding
This work was partially supported by the Brown Institute for Media Innovation, by Google through their affiliation with Stanford Institute for Human-centered Artificial Intelligence (HAI) and by a Hoffman-Yee HAI grant.